OpenStudy (anonymous):

Please help me. This is the only question i need help on. Ill fan or Medal. 18. A diameter of a circle has endpoints P(–10, –2) and Q(4, 6). Find the center of the circle. Find the radius. If your answer is not an integer, express it in radical form. Write an equation for the circle. (3 points)

OpenStudy (anonymous):

center is between P Q

OpenStudy (anonymous):

I think there wanting actual points. For the center i got -3, -2 But i think that is wrong

OpenStudy (anonymous):

O(xp+xq/2,yp+yq/2/)

OpenStudy (anonymous):

your equation confuses me.

OpenStudy (anonymous):

no it is (-3,2)

OpenStudy (anonymous):

-10+4/2=-3 -2+6/2=2

OpenStudy (anonymous):

Ohh okay that makes since now. What about the radius?

OpenStudy (anonymous):

It asks for in in an integer or radical form, which confused me

OpenStudy (anonymous):

PQ=sqrt[(xp-xo)2+(yp-yo)2]

OpenStudy (anonymous):

would it be sqrt85?

OpenStudy (anonymous):

R=PQ/2

OpenStudy (anonymous):

PQ=sqrt(-10-4)2+(-2-6)2=sqrt(14^2+8^2)=sqrt(260) R=sqrt(65)

OpenStudy (anonymous):

approximate R=8.06

OpenStudy (anonymous):

so it would be 65?

OpenStudy (anonymous):

and is that in an interger or radical form? Sorry i know i seem like i dont get any of this. Ive had some things interfere with school this semester.

OpenStudy (anonymous):

sqrt(65) yes

OpenStudy (anonymous):

Thank you. what about the equation? I got (x +3)^2 + (y -2)^2

OpenStudy (anonymous):

which would = 65

OpenStudy (anonymous):

(x +3)^2 + (y -2)^2=65 yes

OpenStudy (anonymous):

Thank you very much. I have 2 questions i couldnt figure out, do you have time to help me?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

write in new post

OpenStudy (anonymous):

okay thanks i did